Assume a hypothetical consumer good X and good Y. the price of good X is 1 and price of good Y is 3 and the consumer budget is birr 10 for the two goods. where: Qx is quantity of good X, Qy is quantity of good Y and TUx and TUy is total Utility from consuming good X and Good Y respectively
Qx. TUx. Qy TUy
0. 0. 0. 0
1. 10 1. 24
2. 19. 2. 45
3. 27 3. 63
4. 34. 4. 78
5. 40. 5. 87
6. 44. 6. 90
Based on this
At what amount of consumption dose diminishing marginal utility startes to occurs for the two goods
1
Expert's answer
2019-10-10T09:26:24-0400

Solution: Find marginal utility for each unit of good

Qx TUx MUx Qy TUy MUy

0 0 0 0 0 0

1 10 10 1 24 24

2 19 9 2 45 23

3 27 8 3 63 18

4 34 7 4 78 15

5 40 6 5 87 9

6 44 4 6 90 3


The price of good X is 1 and price of good Y is 3 and the consumer budget is birr 10 for the two goods, so if I=10, then I=PxQx+PyQy, I=Qx+3Qy.

Qx+3Qy=10.

1)   Qx=1, Qy=3, MU=10+18=28;

2)   Qx=4, Qy=2, MU=7+23=30.


There are no other options for the full use of the budget.

Ansver:  Qx=4, Qy=2.